OpenStudy (anonymous):

Please check my answer on point slopes.. I must have made a mistake! Here is the problem: Find the equation of a line given the point and slope below. Arrange your answer in the form y = mx + b, where b is the constant. (1, 3) m = 2 y = 3x + b ?? Can't I eliminate the b?

OpenStudy (anonymous):

How do I eliminate the b?

jimthompson5910 (jim_thompson5910):

The given point (1,3) is in the form (x,y) So (x,y) = (1,3) which means that x = 1 and y = 3

jimthompson5910 (jim_thompson5910):

So y = 3x + b 3 = 3(1) + b and you can now solve for b

OpenStudy (anonymous):

Okay, b is zero then.

OpenStudy (anonymous):

Except that it seems that the slope is given as 2

OpenStudy (anonymous):

y = 3

jimthompson5910 (jim_thompson5910):

oh glossed over that part, my bad

OpenStudy (anonymous):

So it should be y = 2x + b because m = 2

jimthompson5910 (jim_thompson5910):

y = mx + b y = 2x + b 3 = 2(1) + b now you can solve for b

OpenStudy (anonymous):

y = 2x + 1

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

yes roko

OpenStudy (anonymous):

Thanks.. I wish I could give "best response to both of you!